# From Imitation to Refinement – Residual RL for Precise Assembly

### v3

@@ IV-C Real-World Application
Figure 11: (A) When changing the part appearances from white to black, the policy trained on real data only seizes to function. The behavior becomes erratic, and all trials ended with the robot hitting a velocity limit. (B) When mixing in synthetic data with more diverse colors (see Fig. 9), the policy regains the ability to complete the task, though with lower performance than for white parts.
@@ V Related Works
Training diffusion models with reinforcement learning
The approach in [46, 56] studied how to cast diffusion de-noising as a Markov Decision Process, enabling preference-aligned image generation with policy gradient RL. However, this method is unstable and difficult to tune. [47] introduced a framework for direct diffusion policy fine-tuning, which avoids this instability, but the method is architecture-specific and does not introduce closed-loop control. Other ways to combine diffusion architectures with RL include Q-function-based importance sampling [58], advantage weighted regression [59], or changing the objective into a supervised learning problem with return conditioning [60, 17, 16]. Some have also explored augmenting the de-noising training objective with a Q-function maximization objective [61] and iteratively changing the dataset with Q-functions [75]. Recent work developed techniques for training diffusion policies from scratch [57], leveraging unsupervised clustering and Q-learning ensembles to encourage multi-modal behavior discovery. Our method avoids such complexity involved with directly optimizing diffusion models by using standard PPO to train simple residual policies that correct for errors made by the base policy.
Residual learning in robotics
@@
Learning corrective residual components in conjunction with learned or non-learned “base” models has been widely successful in robotics. Common frameworks include learning residual policies that correct for errors made by a nominal behavior policy [48, 64, 76, 77, 78, 79, 80, 81] and combining learned components to correct for inaccuracies in analytical models for physical dynamics [82, 83, 84] or sensor observations [85]. Residual policies have been used in insertion applications [86], and recent work has applied residual policy learning to the same FurnitureBench task suite we study in this paper [87]. Their approach uses the residual component to model online human-provided corrections via supervised learning, whereas we train our residual policy from scratch with RL using task rewards in simulation.
VI Limitations and Conclusion
Limitations
@@ VI Limitations and Conclusion
The local nature of our residual policies is not well-suited for learning the macro-level behaviors required to recover from large-scale deviations like dropped parts. Our proposed method also struggles in regimes with very high initial scene randomness, as both the base policies and actions produced via RL exploration struggle to deal with out-of-support initial part poses. Furthermore, despite showcasing the advantage of incorporating simulation data, sim-to-real for RGB policies still presents a challenge. There remains a performance gap in both teacher-student distillation and sim-to-real distribution shifts. Future investigations may include better sim-to-real transfer techniques, exploration mechanisms for discovering how to correct large-scale execution errors, tractable interactive learning for real-world policy distillation, and incorporating inductive biases that help generalize to much broader initial state distributions.
Conclusion
This work presents ResiP, a simple yet effective approach for improving behavior-cloned policies on precise manipulation tasks using sparse rewards. Our key insight is recognizing modern chunked BC policies as trajectory planners that require closed-loop corrections for reliable execution. By training a residual policy with RL to provide these corrections, we sidestep the complexities of directly fine-tuning architectures like diffusion models while addressing fundamental limitations of BC – distribution shift and lack of reactivity.
Our results demonstrate that this approach significantly outperforms pure BC methods and direct RL fine-tuning across multiple challenging assembly tasks, improving success rates from $\sim$50-80to over 95% in many cases. Through careful ablations, we show the importance of the fully closed-loop nature of the residual and the choice of a robust base policy, particularly for high-precision tasks like sub-millimeter insertions.
While we demonstrate successful real-world transfer through teacher-student distillation and sim-to-real co-training, the remaining performance gap suggests opportunities for improved sim-to-real procedures through direct vision-based RL or interactive distillation. Additionally, our results with different base architectures suggest that this residual learning approach is valuable for fine-tuning foundation models while preserving diverse capabilities.
@@ Acknowledgments
This work was partly supported by the Sony Research Award and the US Government. The computations in this paper were run on the FASRC cluster, supported by the FAS Division of Science Research Computing Group at Harvard University, and on the MIT Supercloud[88]. Experiment tracking and model checkpoint storage were provided by Weights and Biases.
@@
PA advised the project and provided valuable feedback on project framing and contributions.
@@ A-A Training Hyperparameters
State-based behavior cloning
@@
State-based reinforcement learning
@@
Image-based real-world distillation
For the real-world experiments, we use a separate set of hyperparameters, presented in Tab. X. The main difference is that we found in experimentation that the transformer backbone in  [18] worked better than the UNet for real-world experiments. These models are also operating from RGB observations instead of privileged states, and we provide parameters for the image augmentations applied to the front camera in Tab. XI and the wrist camera in Tab. XII.
@@ B-A Tasks details and reward signal
Furniture assembly tasks
We detail a handful of differentiating properties for each of the three tasks we use in Tab. XIII. one_leg involves assembling 2 parts, the tabletop and one of the 4 table legs. The assembly is characterized as successful if the relative poses between the parts are close to a predefined assembled relative pose. When this pose is achieved, the environment returns a reward of 1. That is, for the one_leg task, the policy received a reward of 1 only at the very end of the episode. For round_table and lamp, which consists of assembling 3 parts together, the policy receives a reward signal of 1 for each pair of assembled parts. E.g. for the lamp task, when the bulb is fully screwed into the base, the first reward of 1 is received, and the second is received when the shade is correctly placed.
Real-to-sim task: mug-rack
@@
High-precision, Factory task: peg-in-hole
@@
Bimanual, high-precision task: biman-insert
@@ Appendix D Visualization of Residual Policy Actions
Appendix E Real-World Results
E-A Quantitative Results Failure Mode Breakdown
@@
E-B Extension of Pipeline to Unseen Settings
@@ Appendix G Extended Limitations and Further Work
Appendix H Further Analysis of Offline versus Online Learning
H-A Distillation scaling analysis
@@
H-B Interactive distillation with DAgger
@@
Appendix I Residual RL ablations
I-A Effect of fully versus partially closed-loop policies
@@
I-B Residual base policy ablation
@@
I-C Residual action scaling parameter ablation

### v4

@@ IV-C Real-World Deployment
Figure 10: (A) When changing the part appearances from white to black, the policy trained on real data only (Real-Only) seizes to function. The behavior becomes erratic, and all trials ended with the robot hitting a velocity limit. (B) When mixing in synthetic data with more diverse colors (see Fig. 3), the policy (Real+Sim) regains the ability to complete the task, though with lower performance than for white parts.
IV-C1 Real-World Performance
In our sim-to-real experiments, following the protocol described in Sec. III-C, we find that the Real+Sim policies achieve significantly higher success rates (50-60%) compared to Real-Only baselines (20-30%) on the one_leg task (Tab. II). Qualitatively, the co-trained policy (Real+Sim) exhibits smoother behavior with fewer erratic movements that exceed robot hardware limits. Fig. 9 shows typical behaviors: Row (A) shows successful assembly sequences, while Row (B) demonstrates the primary failure mode—imprecise alignment before release, mirroring challenges observed in simulation.
To assess robustness to visual variations, we tested changing part colors from the white used in data collection to an unseen black. Unsurprisingly, the Real-Only policy fails catastrophically, triggering velocity limits on every trial, as shown in Fig. 10 (A). In contrast, Real+Sim, trained with the same real-world data but with color-randomized synthetic data (see Fig. 3; bottom), maintains basic functionality though with reduced performance compared to the original color scheme. While highlighting the benefits of synthetic data, this also indicates opportunities for improving sim-to-real transfer. See App. E-B for more detailed analysis.
IV-C2 Understanding Performance Limitations
To understand the gap between simulation and real-world performance, we hypothesize three potential limiting factors: (1) the change from state- to vision-based observations, (2) the sim-to-real gap, and (3) the policy distillation process. Comparative experiments between image and state-based students show similar performance gaps from the teacher’s 98% success rate (Fig. 11), ruling out the change in observation modality as the primary bottleneck. Furthermore, our scaling analysis shows that even increasing synthetic trajectories from 10k to 100k only marginally improves success rates from 78% to 80% (LABEL:fig:overview; Right), indicating a fundamental limitation in the policy distillation process itself rather than purely sim-to-real challenges. These findings motivate future investigations into better approaches for vision-based policy distillation, particularly focusing on online learning and closed-loop control. See App. E-A for detailed analysis.
Figure 11: A comparison of performance between state and vision policies in policy distillation. The Diffusion Policy (DP) baseline achieves similar performance with both state and vision inputs when trained directly on the same 50 demonstrations (left), indicating that the vision modality does not inherently limit performance. However, we observe a significant performance gap when distilling our state-based ResiP policy into a vision-based policy (right). This consistent performance gap suggests the challenge lies in the distillation process rather than the vision modality, as direct vision-based training shows no such degradation.
@@ V Related Works
V-A Training diffusion models with reinforcement learning
A fundamental challenge in applying RL to diffusion models is that the final action probabilities are not directly accessible due to the iterative nature of the denoising process, making policy gradient methods difficult to apply. Recent work has explored various approaches to combining diffusion models with RL [73, 74, 40, 75, 41]. Some approaches cast diffusion de-noising as a Markov Decision Process [40, 74], enabling preference-aligned image generation with policy gradient RL, but suffer from training instability. While [41] introduced more stable direct diffusion policy fine-tuning, their method remains architecture-specific and lacks closed-loop control.
Other approaches include Q-function-based importance sampling [65], advantage weighted regression [76], and return-conditioned supervised learning [77, 18, 17]. Some methods augment the denoising objective with Q-function maximization [78] or iteratively update the dataset using Q-functions [79]. However, these approaches mainly enable better extraction or stitching of existing behaviors rather than learning new, corrective behaviors. Recent work on training diffusion policies from scratch [75] uses complex mechanisms like unsupervised clustering and Q-learning ensembles for multi-modal behavior discovery.
Beyond the iterative denoising challenge, modern BC approaches introduce additional complications through action chunking, where policies predict sequences of multiple future actions. While this improves BC performance, it creates significant challenges for RL fine-tuning by expanding the action space—for instance, chunks of 8 actions result in an 8-fold increase in action dimensionality. This issue affects most modern BC architectures like ACT [6, 15], as they rely on action chunking. Policy gradient methods struggle with such high-dimensional action spaces, particularly when applied to deep neural networks [80]. Furthermore, recent work shows that RL fine-tuning of large pre-trained models can lead to forgetting of pre-training capabilities [81]. Our method avoids these problems by keeping the base policy frozen and training only a small residual model, which preserves the pre-trained capabilities and enables stable policy gradient training with closed-loop control.
V-B Residual learning in robotics
@@
Learning corrective residual components in conjunction with learned or non-learned “base” models has been widely successful in robotics. Common frameworks include learning residual policies that correct for errors made by a nominal behavior policy [42, 43, 44, 45, 46, 47, 48, 49] and combining learned components to correct for inaccuracies in analytical models for physical dynamics [82, 83, 84] or sensor observations [85]. Unlike prior residual policy learning approaches, our method uniquely combines RL-based training with running the base at a lower frequency, with the residual providing corrections at every control step.
Residual policies have been used in insertion applications [86], and recent work has applied residual policy learning to the same FurnitureBench task suite we study in this paper [87]. Their approach uses the residual component to model online human-provided corrections via supervised learning, whereas we train our residual policy from scratch with RL using task rewards in simulation.
VI Discussion
@@
The local nature of our residual policies is designed to complement rather than replace the trajectory planning capabilities of the base policy. While this design enables precise corrections, our approach still relies on the base policy for macro-level behaviors. As such, our proposed method struggles in regimes with very high initial scene randomness, as both the base policies and actions produced via RL exploration struggle to deal with out-of-support initial part poses. Our imitation learning scaling analyses were conducted using a dataset from an RL expert, not human demonstrations. These two demonstration sources will likely have different distributions, which may change the analyses. However, acquiring 100k demonstrations from a human demonstrator was not feasible in the present work. Furthermore, despite showcasing the advantage of incorporating simulation data, sim-to-real for vision-based policies still presents a challenge. There remains a performance gap in both teacher-student distillation and sim-to-real distribution shifts. Future investigations may include better sim-to-real transfer techniques, exploration mechanisms for discovering how to correct large-scale execution errors, tractable interactive learning for real-world policy distillation, and incorporating inductive biases (like [88]) that help generalize to broader initial state distributions.
@@ Acknowledgments
This work was partly supported by the Sony Research Award, the US Government, and the Hyundai Motor Company. The computations in this paper were run on the FASRC cluster, supported by the FAS Division of Science Research Computing Group at Harvard University, and on the MIT Supercloud[89]. Experiment tracking and model checkpoint storage were provided by Weights and Biases.
@@
PA advised the project and provided valuable feedback on project framing, contributions, and writing.
@@ A-A Training Hyperparameters
A-A1 State-based behavior cloning
@@
A-A2 State-based reinforcement learning
@@
A-A3 Image-based real-world distillation
We use a separate set of hyperparameters for real-world experiments, presented in Tab. X. The main difference is that we found in experimentation that the transformer backbone in  [19] worked better than the UNet for real-world experiments. These models are also operating from RGB observations instead of privileged states, and we provide parameters for the image augmentations applied to the front camera in Tab. XI and the wrist camera in Tab. XII.
@@ B-A Tasks details and reward signal
B-A1 Furniture assembly tasks
We detail a handful of differentiating properties for each of the three tasks we use in Tab. XIII. one_leg involves assembling 2 parts, the tabletop and one of the 4 table legs. The assembly is successful if the relative poses between the parts are close to a predefined assembled relative pose. When this pose is achieved, the environment returns a reward of 1. That is, for the one_leg task, the policy received a reward of 1 only at the very end of the episode. For round_table and lamp, which consists of assembling 3 parts together, the policy receives a reward signal of 1 for each pair of assembled parts. E.g. for the lamp task, when the bulb is fully screwed into the base, the first reward of 1 is received, and the second is received when the shade is correctly placed.
B-A2 Real-to-sim task: mug-rack
@@
B-A3 High-precision, Factory task: peg-in-hole
@@
B-A4 Bimanual, high-precision task: biman-insert
@@ Appendix D Visualization of Residual Policy Actions
Appendix E Extended Vision-Based Results and Analyses
E-A Performance Impact of Distillation
Next, we study how the quantity and quality of data generated by a ResiP policy impact the performance of vision-based student policies in real-world evaluations. We generate this data by collecting successful trajectories from the ResiP teacher across varied initial states and rendering corresponding camera observations. A vision-based student policy—which shares the same architecture as the teacher but includes an additional image encoder—distilled from $\sim$1,000 teacher trajectories reached 73% success on one_leg, outperforming the 50% achieved by training the vision policy directly on human demos (see Fig. 11). However, we observe a performance gap between the RL-trained ResiP teacher (98%) and the distilled vision-based student (73%), even after performance saturates with additional data. To investigate whether this gap stems from the change to visual input, we compared distillation performance between image-based and state-based students using the same number of trajectories. Their comparable performance suggests that the modality shift is not the primary cause of the performance gap. While DAgger-style online distillation might improve performance, we focused on offline distillation as it better reflects real-world deployment constraints.
Therefore, we examine the impact of the distillation dataset size. Here, we scale up the number of state-based rollouts from the trained RL policy and distill these to a state-based student. In LABEL:fig:overview (Right), we observe that performance increases with more data, from 78% success rate at 10k trajectories to 80% at 100k trajectories, though not reaching the teacher policy’s 98% success rate. The same trend is evident in Fig. 5 (though the saturation occurs earlier). These results demonstrate how simulation-based distillation can complement existing training approaches by enabling the rapid generation of large-scale synthetic datasets at a minimal cost. Beyond the data volume advantages, our RL teacher exhibits qualitatively different behaviors, such as faster movements and improved corrective actions, suggesting that this synthetic data captures valuable task strategies that could be expensive or impractical to demonstrate manually.
While DAgger [1] demonstrates strong sample efficiency - achieving better performance than BC with just $\sim$10k gradient steps (800 rollouts)—it requires an expert policy for online data collection. Although we could theoretically apply DAgger in simulation using our RL-trained expert policy (as demonstrated by [16] for point cloud inputs), this introduces significant complexity to the pipeline and its effectiveness for RGB image-based distillation remains an open question for future work. Instead, we demonstrate that a simple approach combining offline rendering of synthetic trajectories with real-world co-training can achieve reasonable performance.
E-B Real-World Evaluation
Finally, we evaluate the real-world performance of a sim-to-real policy trained on a mixture of a few (10/40) real-world demonstrations (Real+Sim) and simulation data generated by the trained residual RL policy. We compare the co-trained policy to a baseline model trained only on real-world demonstrations (Real-Only). We compare the success rates achieved by each policy on two sets of 10 trials for the one_leg task. In the first set, we randomize part poses, while in the second set, we randomize obstacle poses (i.e., insertion location in the workspace).
We compare the co-trained policy to a baseline model trained only on real-world demonstrations. We define an evaluation grid spanning the same ranges as the low randomization setting from the FurnitureBench simulation environment. We evaluate each policy on two sets of 10 trials for the one_leg task, with grid points sampling either part poses or obstacle poses (i.e., insertion location in the workspace). Each method is evaluated on the same set of grid positions to ensure fair comparison.
The results in Tab. II show that incorporating simulation data improves real-world performance (e.g., increasing task completion rate from 20-30% to 50-60%). Qualitatively, the sim-to-real policy exhibits smoother behavior and makes fewer erratic movements that might exceed the robot’s physical limits. Fig. 9 illustrates this through example trajectories: Row (A) shows successful executions where the robot completes the full assembly sequence, while Row (B) demonstrates the most common failure mode where, despite successfully grasping and transporting the parts, the policy fails to achieve precise alignment between the table leg and the hole before releasing. This misalignment failure pattern mirrors what we observe in the simulation, suggesting consistent challenges in achieving the required precision for insertion tasks.
To further probe the robustness conferred by training in simulation, we created a task variation where the part colors are changed from black to white. When rolling out the policy trained on real demos of white parts, Real-Only, the robot exhibited erratic behavior that caused the hardware to reach velocity limits on every trial we ran, as shown in Fig. 10 (A). When including synthetic data rendered with parts in black, the resulting policy (Real+Sim-DR) can perform the task again (see Fig. 10). The resulting performance was still inferior to the performance on white parts, which motivates further work on closing the sim-to-real gap.
E-C Quantitative Results Failure Mode Breakdown
@@
E-D Extension of Pipeline to Unseen Settings
@@ Appendix G Extended Limitations and Further Work
Appendix H Why Action Chunking and Diffusion Policies?
Appendix I Further Analysis of Offline versus Online Learning
I-A Distillation scaling analysis
@@
I-B Interactive distillation with DAgger
@@
Appendix J Residual RL ablations
J-A Effect of fully versus partially closed-loop policies
@@
J-B Residual base policy ablation
@@
J-C Residual action scaling parameter ablation